Question 1

Fixed cost per unit continuously declines as production increases (within the relevant range). This is called economies of scale.

True

False

Question 2

The coefficient of determination or R squared, in regression analysis, should be less than 0.75 to indicate a strong cost equation.

True

False

Question 3

ABC Co. has fixed costs of $85,000. ABC wants to achieve an after-tax profit of $10,000. The selling price per unit is $100, and the variable cost per unit is $25. The company has a tax rate of 20%. How many units must ABC sell to acheive its after tax target profit goal?

Question 4

Advertising represents a variable cost.

True

False

Question 5

The following is historical costs and usage for electricity costs:

                      Electricity Cost             Kilowatt Hours

January              $1500                        2600

February            $1200                        1800

Using the high/low method, what is the projected fixed cost? Omit dollar signs and commas!

Question 6

Mixed costs have fixed and variable components.

True

False

Question 7

The relevant range represents where all cost assumptions remain valid.

True

False

Question 8

XYZ Company has fixed costs of $100,000. The selling price per unit is $80 and the variable cost per unit is $40. What is the break even sales dollars?

Omit dollar signs and commas!

Question 9

A step fixed costs represents a fixed cost that varies in direct proportion to the activity level

True

False

Question 10

Total fixed costs change in direct proportion to the level of production.

True

False
Answer #1

Answers

  • Question 1: Fixed cost per unit continuously declines as production increases (within the relevant range). This is called economies of scale.

Answer = TRUE; Fixed cost remains same is totality but changes per unit basis.

  • Question 2: The coefficient of determination or R squared, in regression analysis, should be less than 0.75 to indicate a strong cost equation.

Answer = FALSE, its not essential

  • Question 3: ABC Co. has fixed costs of $85,000. ABC wants to achieve an after-tax profit of $10,000. The selling price per unit is $100, and the variable cost per unit is $25. The company has a tax rate of 20%. How many units must ABC sell to acheive its after tax target profit goal?

A

After Tax profit

$            10,000.00

B = A/(100% - 20%) or A/80%

Before tax profot would be

$            12,500.00

C

Fixed Cost

$            85,000.00

D = B+C

Total Contribution margin required

$            97,500.00

E = 100 - 25

Contribution margin per unit

$                    75.00

F = D/E

No. of units required

                        1,300 = Answer

  • Question 4: Advertising represents a variable cost.
    The answer is FALSE. Advertisement Cost represent FIXED COST
  • Question 5: Answer = $ 525 = Fixed Cost

Question 5

Months

Units

Cost

High Level

Jan

                            2,600

$                     1,500.00

Low Level

Feb

                            1,800

$                     1,200.00

Difference

                                800

$                         300.00

Difference in Cost

$                                                                   300.00

Difference in units

                                                                            800

Variable cost per unit

$                                                                       0.38

Working

High Level

Low Level

A

Total Cost

$                   1,500.00

$                     1,200.00

B

Total Units

2600

1800

C

Variable cost per unit

$                           0.38

$                             0.38

D = B x C

Total Variable cost

$                       975.00

$                         675.00

E = A - D

Total Fixed Cost

$                       525.00

$                         525.00

--Answer: FIXED COST = $ 525.00

  • Question 6: Mixed costs have fixed and variable components.

TRUE, that’s why they are called MIXED cost.

  • Question 7: The relevant range represents where all cost assumptions remain valid.

Answer =True. It assumes that Fixed cost will remain same and variable cost will change with a change in volume.

  • Question 8: XYZ Company has fixed costs of $100,000. The selling price per unit is $80 and the variable cost per unit is $40. What is the break even sales dollars?

Answer: = $ 200,000

A

Fixed Cost

$          100,000.00

B = 80 - 40

Contribution margin per unit

$                    40.00

C = A/B

Break Even in units

                        2,500

D = C x $ 80

Break Even in Dollars

$          200,000.00

  • Question 9: A step fixed costs represents a fixed cost that varies in direct proportion to the activity level.

FALSE. It varies, but not in direct proportion to the activity ;evel.

  • Question 10: Total fixed costs change in direct proportion to the level of production.

FALSE. Variable cost does that, Fixed Cost stays the same.